OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1,3) and (1,11)?

OpenStudy (bahrom7893):

M = (y-y1)/(x-x1)

OpenStudy (anonymous):

y=4x+7

OpenStudy (bahrom7893):

M = (11-3)/(1+1) = 8/2 = 4